Evaluate 1-2cos²25/(1-2sin²65)
1
Expert's answer
2019-04-29T11:05:45-0400

Evaluate 1-2cos²25/(1-2sin²65)



Solution:

"1- (2cos\u2061^225\\circ)\/(1-2sin\u2061^265\\circ)\n=""=1-(2(1+cos\u206150\\circ)\/2)\/cos\u2061130\\circ \n=""=1-(1+cos\u206150\\circ)\/(-cos\u206150\\circ )\n=""=1+(1+cos\u206150\\circ)\/cos\u206150\\circ \n=""=1+1\/cos\u206150\\circ +cos\u206150\\circ\/cos\u206150\\circ\n=""=1+1+1\/cos\u206150\\circ =2+1\/cos\u206150\\circ \u22483.55572"

Answer: 3.55572